OpenStudy (anonymous):

Solve (: t/3 - 1/2 = 1. A.4.1 B. 5.0 C. 4.2 D. 4.5 Note-- I can't reply

OpenStudy (camerondoherty):

D

OpenStudy (anonymous):

4.5

OpenStudy (anonymous):

its like this.. t/3=1+1/2=3/2 =>t=3*3/2=9/2=4.5